OpenStudy (anonymous):

How do you write a equation for the transformation of y=x: with a vertical stretch by a factor of 2

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (jdoe0001):

y = A ( Bx + C ) + D A = vertical stretch/expand by a factor of A

OpenStudy (anonymous):

So it would be ... y= 2( and then im lost.